OpenStudy (anonymous):

solve the equation 4y+5=-31

OpenStudy (anonymous):

-9

OpenStudy (saifoo.khan):

Do you want to graph it? @buddhababy002

OpenStudy (anonymous):

no i dont need graph, just how to solve it.

OpenStudy (anonymous):

I think he has written solve the equation.. @saifoo.khan

OpenStudy (saifoo.khan):

@buddhababy002 , the equation is already solved and it's in standard form.

OpenStudy (anonymous):

Ha ha.. Solve for y @saifoo.khan

OpenStudy (saifoo.khan):

@waterineyes , please let me solve the problem. :)

OpenStudy (anonymous):

Are you finding any x there ??

OpenStudy (anonymous):

Carry on @saifoo.khan

OpenStudy (saifoo.khan):

@buddhababy002 : Solve this equation, make y the subject. This will help you to convert the line in slope-intercept form.

OpenStudy (anonymous):

you do not need to, just move 5 to the right side, and divide both sides by 4

OpenStudy (saifoo.khan):

lol, really?

OpenStudy (anonymous):

Subtract 5 both the sides: \[4y = -31 - 5 \implies 4y = -36\] Divide by 4 now: \[y = \frac{-36}{4} \implies y = -9\]

OpenStudy (saifoo.khan):

@waterineyes : haha. ;)

OpenStudy (anonymous):

@saifoo.khan I think you are using your brain more than required.. Ha ha ha..

OpenStudy (anonymous):

so u subtract both sides by 5 cause it doesn't have the variable right?

OpenStudy (anonymous):

I subtract 5 because we want to make variable on one side and constant terms on other side..

OpenStudy (saifoo.khan):

@waterineyes , hahah right! im used to that "more brain using" method actually. LOLOLOL

OpenStudy (anonymous):

ah i see.. okay thank you.

OpenStudy (anonymous):

Welcome..

OpenStudy (saifoo.khan):

nvm.